OpenStudy (anonymous):

X squared minus 9 all over x squared plus 3 x simplified

OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

y = (x^2 - 3)/(x^2 + 3x) = (x -3)(x + 3)/x(x + 3) = (x - 3)/x